# Knots and variance ordering of sequential Monte Carlo algorithms

### v1

@@ 1 Introduction
In the canonical SMC algorithm, the bootstrap particle filter (Gordon et al., 1993), information is incorporated into the particle system according to the time of observation. Methods such as the auxiliary particle filter (Pitt and Shephard, 1999; Johansen and Doucet, 2008), look-ahead particle filters (Lin et al., 2013), and model twisting methods (Guarniero et al., 2017; Heng et al., 2020), define new particle filters that incorporate varying degrees of future information into the current time step. Idealised versions of these algorithms reduce or eliminate variance in the particle system, producing more accurate particle approximations at the cost of increased computation. Typically, these idealised filters are not computationally tractable or are prohibitively expensive to calculate, so practical methods frequently employ approximations to this ideal filter.
@@
This paper contributes knots for Feynman–Kac models, a technique to reduce the asymptotic variance for particle approximations. We generalise and unify ‘full’ adaptation and Rao–Blackwellisation, improving the theoretical underpinning for these methods, whilst providing a highly flexible strategy for designing new algorithms with guaranteed asymptotic variance reduction. Further, we resolve the discrepancy between the theoretical analysis of full-adapted auxiliary particle filters and their demonstrated empirical performance. We demonstrate that a small change to particle filters with ‘full’ adaptation can guarantee an asymptotic variance reduction, resolving the counter-example in Johansen and Doucet (2008).
@@
The paper is structured as follows. We first review the background of SMC algorithms, including Feynman–Kac models and the asymptotic variance of particle approximations in Section 2. Knots are introduced in Section 3 where we discuss their properties as well as invariances and equivalences of models before and after the application of a knot. Section 4 contains our main variance reduction results for knots, whereas Section 5 discusses the optimality of knots. Section 6 provides variance and optimality results for terminal knots which require special treatment. Lastly, Section 7 contains examples including ‘full’ adaptation and Rao–Blackwellisation as special cases of knots, and an illustrative example that is a hybrid between these two cases.
@@ 2.1 Notation
Let $(\mathsf{X},\mathcal{X})$ be a measurable space. If $\mu$ is a measure on $(\mathsf{X},\mathcal{X})$ and the function $\varphi:\mathsf{X}\rightarrow\mathbb{R}$} then let $\mu(\varphi)=\int\varphi(x)\mu(\mathrm{d}x)$ and if $S\in\mathcal{X}$ then let $\mu(S)=\mu(1_{S})$. A degenerate probability measure at $x\in\mathsf{X}$ is denoted by $\delta_{x}$. We use $\mathcal{L}(\mu)$ to denote the class of functions that are $L^{1}$-integrable with respect to a measure $\mu$.
@@
We make use of twisted Markov kernels and will use the following superscript notation when describing these. If $K:(\mathsf{X},\mathcal{Y})\rightarrow[0,1]$ is a Markov kernel and $H:\mathsf{Y}\rightarrow[0,\infty)$ is $L^{1}$-integrable w.r.t. $K(x,\cdot)$ for all $x\in\mathsf{X}$, let $K^{H}:(\mathsf{X},\mathcal{Y})\rightarrow[0,1]$ be defined such that
$K^{H}(x,\mathrm{d}y)=\begin{cases}\frac{K(x,\mathrm{d}y)H(y)}{K(H)(x)}&~\text{if}~K(H)(x)>0,\\ Q(x,\mathrm{d}y)&~\text{if}~K(H)(x)=0,\end{cases}$
for an arbitrary Markov kernel $Q:(\mathsf{X},\mathcal{Y})\rightarrow[0,1]$. The choice $Q$ is arbitrary in our context as this state of the Markov kernel will always be zero weighted. Its specification can be safely ignored in particle filter implementations but is useful for proofs. Similarly, if $\mu$ is a probability measure on $(\mathsf{Y},\mathcal{Y})$ then $\mu^{H}$ will be defined as $\mu^{H}(\mathrm{d}y)=\frac{\mu(\mathrm{d}y)H(y)}{\mu(H)}$ if $\mu(H)>0$ and $\mu^{H}(\mathrm{d}y)=P(\mathrm{d}y)$ if $\mu(H)=0$ for an arbitrary probability distribution $P$ on $(\mathsf{Y},\mathcal{Y})$. If $K$ is the identity kernel or degenerate probability measure, we take $K^{H}=K$ by convention.
@@ 2.2 Discrete-time Feynman–Kac models
Consider measurable spaces $(\mathsf{X}_{p},\mathcal{X}_{p})$ for $p\in[0\,{:}\,n]$, then let $M_{0}:\mathcal{X}_{0}\rightarrow[0,1]$ be a probability measure, $M_{p}:(\mathsf{X}_{p-1},\mathcal{X}_{p})\rightarrow[0,1]$ for $p\in[1\,{:}\,n]$ be Markov kernels, and consider potential functions $G_{p}:\mathsf{X}_{p}\rightarrow[0,\infty)$ that are $L^{1}$-integrable with respect to $M_{p}(x,\cdot)$ for all $x\in\mathsf{X}_{p-1}$ and $p\in[0\,{:}\,n]$.
@@ 2.4 Asymptotic variance of particle approximations
for a particle filter following Algorithm 1. Under various conditions, particle approximations relate to this variance by way of Central Limit Theorems (CLT, Del Moral, 2004). For the terminal predictive (probability) measures we have $\gamma_{n}^{N}(\varphi)\rightarrow\gamma_{n}(\varphi)$} and $\eta_{n}^{N}(\varphi)\rightarrow\eta_{n}(\varphi)$} almost surely as $N\rightarrow\infty$ with
$N\,\mathrm{Var}\left\{\gamma_{n}^{N}(\varphi)/\gamma_{n}(1)\right\}\rightarrow\sigma^{2}(\varphi),\quad N\,\mathbb{E}\left[\left\{\eta_{n}^{N}(\varphi)-\eta_{n}(\varphi)\right\}^{2}\right]\rightarrow\sigma^{2}(\varphi-\eta_{n}(\varphi)),$
from Lee and Whiteley (2018) for example. When $\sigma^{2}(\varphi)$ or $\hat{\sigma}^{2}(\varphi)$ are finite, such a CLT is useful to characterise the theoretical performance of a particle filter using the variance term. For general Feynman–Kac models, CLT statements are frequently made under the assumption of bounded potential functions and a bounded test function for clarity, but this need not be the case.
Our analysis only requires that the asymptotic variance exists and is finite. As such, for the predictive measure of a given Feynman–Kac model $\mathcal{M}$, we will consider functions $\varphi\in\mathcal{F}(\mathcal{M})$ such that $\mathcal{F}(\mathcal{M})=\{\varphi\in\mathcal{L}(\eta_{n}):\sigma^{2}(\varphi)<\infty\}$. Analogous CLTs also hold for the updated marginal (probability) measures by using $\hat{\sigma}^{2}(\varphi)$ and $\hat{\sigma}^{2}(\varphi-\hat{\eta}_{n}(\varphi))$ in place of $\sigma^{2}(\varphi)$ and $\sigma^{2}(\varphi-\eta_{n}(\varphi))$ respectively, where
@@ 3 Tying knots in Feynman–Kac models
We will introduce knots precisely in Section 3.1 and a convenient notion for the simultaneous application of knots, knotsets in Section 3.2. Knots at time $n$ are considered later, in Section 6, as terminal knots require special treatment and have a smaller scope compared to regular knots (time $t<n$).
@@ 3.1 Knots
The knot operator preserves the terminal predictive and updated measures of the Feynman–Kac model, the predictive and updated path measures (see Proposition 3.13). Besides preserving key measures, the knot operator preserves the horizon of the Feynman–Kac model it is applied to, which is crucial for our comparisons of the asymptotic variance of particle estimates with and without knots.
@@
Trivial knots do not change how the information at time $t$ (the potential) is incorporated into the Feynman–Kac model and do not change the asymptotic variance. On the other hand, we can define an adapted knot which fully adapts $M_{t}$ to the information at time $t$. In fact, any knot can be thought of living on a spectrum between a trivial knot and an adapted knot. We discuss the optimality of adapted knots in Section 5.
@@ 3.2 Knotsets
A knot is the elementary operator we consider, in the sense that it is the minimal modification of a Feynman–Kac model for which we can prove a variance reduction. However, it is natural for a horizon $n$ model to consider a set of knots acting on many time points. Knots can be applied sequentially, but it is convenient to consider a set of knots that can be applied simultaneously. As such, we will now define a generalisation of knots and their associated operator, the knotset and knotset operator.
@@
A knotset $\mathcal{K}=(R_{0:n-1},K_{0:n-1})$ is specified by $n$ knots of the form $\mathcal{K}_{p}=(p,R_{p},K_{p})$ for $p\in[0\,{:}\,n-1]$. Such a knotset is compatible with $\mathcal{M}\in\mathscr{M}_{n}$ if every $(p,R_{p},K_{p})$-knot is compatible with $\mathcal{M}$ for $p\in[0\,{:}\,n-1]$.
@@
Let $\mathcal{K}=(R_{0:n-1},K_{0:n-1})$ be a knotset compatible with $\mathcal{M}$. The knotset operation is defined as $\mathcal{K}\ast\mathcal{M}=\mathcal{K}_{0}\ast\mathcal{K}_{1}\ast\cdots\ast\mathcal{K}_{n-1}\ast\mathcal{M}$ where $\mathcal{K}_{p}=(p,R_{p},K_{p})$ for $p\in[0\,{:}\,n-1]$.
@@
We refer to $\mathcal{M}^{\ast}=\mathcal{K}\ast\mathcal{M}$ for a knotset $\mathcal{K}$ as a knot-model and provide the form of $\mathcal{M}^{\ast}$ in Proposition 3.9. The proof is trivial due to the descending order of knot applications specified in Definition 3.8. The knotset operator also inherits right-associativity from knots.
@@
Consider a knotset $\mathcal{K}=(M_{0:n-1},K_{0:n-1})$ where $K_{p}=\mathrm{Id}$ for $p\in[0\,{:}\,n-1]$ applied to model $\mathcal{M}=(M_{0:n},G_{0:n})$. The resulting model $\mathcal{K}\ast\mathcal{M}=\mathcal{M}$.
@@
Though knotsets can change the Feynman–Kac model they are applied to, some quantities remain unchanged whilst others can be expressed in terms of the original model. We note these invariances and equivalences now.
@@
Proof.
Then by Proposition A.1, $\gamma_{p-1}^{\ast}\{K_{p-1}(G_{p-1})\cdot K_{p-1}^{G_{p-1}}R_{p}(\varphi)\}=\gamma_{p-1}^{\ast}K_{p-1}\{G_{p-1}\cdot R_{p}(\varphi)\}$, and hence for $p\in[1\,{:}\,n-1]$,
@@
Proof.
For Part 1, we expand (1) at time $n$ to get
$=\hat{\gamma}_{p-2}R_{n-1}K_{n-1}[G_{n-1}\cdot M_{n}(\varphi)]$
$=\hat{\gamma}_{p-2}M_{n-1}[G_{n-1}\cdot M_{n}(\varphi)]$
$=\gamma_{n}(\varphi)$
from Proposition 3.9, Proposition 3.12, and Proposition A.1. Since $G_{n}^{\ast}=G_{n}$ the updated terminal measures are also equivalent.
Part 2 follows directly from Part 1 by normalising.
@@ 4 Variance reduction and ordering from knots
Our main result is given in Theorem 4.1 where we state that applying knots to a Feynman–Kac model reduces the variance of particle approximations for all relevant functions. Having already established the equivalence of all terminal marginal (probability) measures in Proposition 3.13, we proceed to considering the asymptotic variances of particle approximations to these quantities. We denote the (probability) measures of the relevant knot-model $\mathcal{M}^{\ast}$ as $\gamma_{n}^{\ast}$, $\hat{\gamma}_{n}^{\ast}$, $\eta_{n}^{\ast}$, $\hat{\eta}_{n}^{\ast}$ and use $\hat{\sigma}^{2}_{\ast}$ and $\hat{\sigma}^{2}_{\ast}$ for the predictive and updated asymptotic variance respectively.
@@
Proof.
From Proposition A.2 we have $Q_{p,n}^{\ast}=K_{p}Q_{p,n}$ for $p\in[0\,{:}\,n-1]$, and from Proposition 3.12 $\gamma_{0}^{\ast}=R_{0}$ and $\gamma_{p}^{\ast}=\hat{\gamma}_{p-1}R_{p}$ for $p\in[1\,{:}\,n-1]$. Therefore, using Jensen’s inequality, for $p\in[1\,{:}\,n-1]$
To quantify the reduction in variance, we can see that
$=R_{0}\left[K_{0}\{Q_{0,n}(\varphi)^{2}\}-K_{0}Q_{0,n}(\varphi)^{2}\right]$
$=R_{0}\left[\mathrm{Var}_{K_{0}}\{Q_{0,n}(\varphi)\}\right]$
@@
Note that the variance reduction excludes a contribution from time $n$ due to the absence of a knot at the terminal time. We can also define a terminal time $(n,R,K)$-knot analogously to the knots discussed thus far. However, such a terminal knot will only guarantee a variance reduction of the normalising constant estimate, $\gamma_{n}(1)$. We introduce and discuss general terminal knots for specific test functions in Section 6.
@@
Proof.
[…]
The differences in the asymptotic variances stated in Corollary 4.2 are straightforward to deriving using the quantitative result in Theorem 4.1, as such we suppress them here.

### v2

@@ 1 Introduction
In the canonical SMC algorithm, the bootstrap particle filter (Gordon et al., 1993), information is incorporated into the particle system according to the time of observation. Methods such as the auxiliary particle filter (Pitt and Shephard, 1999; Johansen and Doucet, 2008), look-ahead particle filters (Lin et al., 2013), and model twisting methods (Guarniero et al., 2017; Heng et al., 2020), define new particle filters that incorporate varying degrees of future information into the current time step. Idealised versions of these algorithms reduce or eliminate variance in the particle system, producing more accurate particle approximations at the cost of increased computation. Typically, these idealised filters are not computationally tractable or are prohibitively expensive to calculate, so practical methods frequently employ approximations to an ideal filter.
@@
This paper contributes knots for Feynman–Kac models, a technique to reduce the asymptotic variance for particle approximations. We generalise and unify ‘full’ adaptation and Rao–Blackwellisation, improving the theoretical underpinning for these methods, whilst providing a highly flexible strategy for designing new algorithms with guaranteed asymptotic variance reduction. Further, we resolve the discrepancy between the theoretical analysis of fully-adapted auxiliary particle filters and their demonstrated empirical performance. We demonstrate that a small change to particle filters with ‘full’ adaptation can guarantee an asymptotic variance reduction, resolving the counterexample in Johansen and Doucet (2008).
@@
The paper is structured as follows. We first review the background of SMC algorithms, including Feynman–Kac models and the asymptotic variance of particle approximations in Section 2. Knots are introduced in Section 3 where we discuss their properties as well as invariances and equivalences of models before and after the application of a knot. Section 4 contains our main variance reduction results for knots, whereas Section 5 discusses the optimality of knots. Section 6 provides variance and optimality results for terminal knots which require special treatment. Lastly, Section 7 contains examples including ‘full’ adaptation and Rao–Blackwellisation as special cases of knots, and an illustrative example that is a hybrid between these two cases. Proofs of results are deferred to the appendix.
@@ 2.1 Notation
Let $(\mathsf{X},\mathcal{X})$ be a measurable space. If $\mu$ is a measure on $(\mathsf{X},\mathcal{X})$ and the function $\varphi:\mathsf{X}\rightarrow\mathbb{R}$} then let $\mu(\varphi)=\int\varphi(x)\mu(\mathrm{d}x)$ and if $S\in\mathcal{X}$ then let $\mu(S)=\mu(1_{S})$. We use $\mathcal{L}(\mu)$ to denote the class of functions that are integrable with respect to a measure $\mu$.
@@
We make use of twisted Markov kernels and will use the following superscript notation when describing these. If $K:(\mathsf{X},\mathcal{Y})\rightarrow[0,1]$ is a Markov kernel and $H:\mathsf{Y}\rightarrow[0,\infty)$ is integrable w.r.t. $K(x,\cdot)$ for all $x\in\mathsf{X}$, let $K^{H}:(\mathsf{X},\mathcal{Y})\rightarrow[0,1]$ be defined such that
$K^{H}(x,\mathrm{d}y)=\begin{cases}\frac{K(x,\mathrm{d}y)H(y)}{K(H)(x)}&~\text{if}~K(H)(x)>0,\\ K(x,\mathrm{d}y)&~\text{if}~K(H)(x)=0.\end{cases}$
The choice to use $K$ when $K(H)(x)=0$ is somewhat arbitrary, but useful for formalising our mathematical arguments. Particle filter approximations do not depend on this choice. Similarly, if $\mu$ is a probability measure on $(\mathsf{Y},\mathcal{Y})$ then $\mu^{H}$ will be defined as $\mu^{H}(\mathrm{d}y)=\frac{\mu(\mathrm{d}y)H(y)}{\mu(H)}$ if $\mu(H)>0$ and $\mu^{H}(\mathrm{d}y)=\mu(\mathrm{d}y)$ if $\mu(H)=0$. If $K$ is the identity kernel or degenerate probability measure, we take $K^{H}=K$ by convention.
@@ 2.2 Discrete-time Feynman–Kac models
Consider measurable spaces $(\mathsf{X}_{p},\mathcal{X}_{p})$ for $p\in[0\,{:}\,n]$, then let $M_{0}:\mathcal{X}_{0}\rightarrow[0,1]$ be a probability measure, $M_{p}:(\mathsf{X}_{p-1},\mathcal{X}_{p})\rightarrow[0,1]$ for $p\in[1\,{:}\,n]$ be Markov kernels, and consider potential functions $G_{p}:\mathsf{X}_{p}\rightarrow[0,\infty)$ that are integrable with respect to $M_{p}(x,\cdot)$ for all $x\in\mathsf{X}_{p-1}$ and $p\in[0\,{:}\,n]$.
@@ 2.4 Asymptotic variance of particle approximations
for a particle filter following Algorithm 1. Under various conditions, particle approximations relate to this variance by way of Central Limit Theorems (CLT, Del Moral, 2004). CLTs establish that the terminal predictive measure, $\gamma_{n}^{N}(\varphi)\rightarrow\gamma_{n}(\varphi)$} in probability as $N\rightarrow\infty$, and
$\sqrt{N}\left(\frac{\gamma_{n}^{N}(\varphi)-\gamma_{n}(\varphi)}{\gamma_{n}(1)}\right)\rightarrow\mathcal{N}\left(0,\sigma^{2}(\varphi)\right),$
in distribution. Whilst for the terminal predictive probability measure, $\eta_{n}^{N}(\varphi)\rightarrow\eta_{n}(\varphi)$} in probability as $N\rightarrow\infty$, and $\sqrt{N}\left(\eta_{n}^{N}(\varphi)-\eta_{n}(\varphi)\right)\rightarrow\mathcal{N}\left(0,\sigma^{2}(\varphi-\eta_{n}(\varphi))\right)$ in distribution. It is also possible to establish convergence in moments (Lee and Whiteley, 2018). In this case, $\gamma_{n}^{N}(\varphi)\rightarrow\gamma_{n}(\varphi)$} and $\eta_{n}^{N}(\varphi)\rightarrow\eta_{n}(\varphi)$} almost surely, whilst
$N\,\mathrm{Var}\left\{\gamma_{n}^{N}(\varphi)/\gamma_{n}(1)\right\}\rightarrow\sigma^{2}(\varphi),\quad N\,\mathbb{E}\left[\left\{\eta_{n}^{N}(\varphi)-\eta_{n}(\varphi)\right\}^{2}\right]\rightarrow\sigma^{2}(\varphi-\eta_{n}(\varphi)).$
When $\sigma^{2}(\varphi)$ is finite, CLTs and convergence statements motivate the use of the asymptotic variance map to characterise the theoretical performance of a particle filter.
We analyse the asymptotic variance map directly, assuming that $\sigma^{2}(\varphi)$ is finite, without additional assumptions. As such, for the predictive measure of a given Feynman–Kac model $\mathcal{M}$, we will consider functions $\varphi\in\mathcal{F}(\mathcal{M})$ such that $\mathcal{F}(\mathcal{M})=\{\varphi\in\mathcal{L}(\eta_{n}):\sigma^{2}(\varphi)<\infty\}$. To interpret our results, we will require a CLT to hold for the model $\mathcal{M}$ in question. For general Feynman–Kac models, for example, bounded potential functions and a bounded test function would suffice, but this need not be the case.
When applicable, analogous CLTs also hold for the updated marginal (probability) measures by using $\hat{\sigma}^{2}(\varphi)$ and $\hat{\sigma}^{2}(\varphi-\hat{\eta}_{n}(\varphi))$ in place of $\sigma^{2}(\varphi)$ and $\sigma^{2}(\varphi-\eta_{n}(\varphi))$ respectively, where
@@ 3 Tying knots in Feynman–Kac models
We will introduce knots formally in Section 3.1 and a convenient notion for their simultaneous application, knotsets, in Section 3.2. Knots at time $n$ are considered later, in Section 6, as terminal knots require special treatment and have a smaller scope compared to regular knots (time $t<n$).
@@ 3.1 Knots
The knot operator preserves the terminal predictive and updated measures of the Feynman–Kac model, the predictive and updated path measures (see Proposition 3.13). Besides preserving key measures, the knot operator preserves the horizon of the Feynman–Kac model it is applied to, which is crucial for our comparisons of the asymptotic variance of particle estimates with and without knots. Figure 1 illustrates the transformation of a Feynman–Kac model to a new model by a knot.
@@
Trivial knots do not change how the information at time $t$ (the potential $G_{t}$) is incorporated into the Feynman–Kac model and do not change the asymptotic variance. On the other hand, we can define an adapted knot which fully adapts $M_{t}$ to the information at time $t$. In fact, any knot can be thought of living on a spectrum between a trivial knot and an adapted knot. We discuss the optimality of adapted knots in Section 5.
@@ 3.2 Knotsets
A knot is the elementary operator we consider, in the sense that it is the minimal modification of a Feynman–Kac model for which we can prove a variance reduction. However, it is natural to consider a set of knots acting on many time points of a model. Further, knots can be applied sequentially, but it is convenient to consider a set of knots that can be applied simultaneously. As such, we will now define a generalisation of knots and their associated operator, the knotset and knotset operator.
@@
A knotset $\mathcal{K}=(R_{0:n-1},K_{0:n-1})$ is specified by $n$ knots of the form $\mathcal{K}_{p}=(p,R_{p},K_{p})$ for $p\in[0\,{:}\,n-1]$. Such a knotset is compatible with $\mathcal{M}\in\mathscr{M}_{n}$ if each $(p,R_{p},K_{p})$-knot is compatible with $\mathcal{M}$ for $p\in[0\,{:}\,n-1]$.
@@
Let $\mathcal{K}=(R_{0:n-1},K_{0:n-1})$ be a knotset compatible with $\mathcal{M}\in\mathscr{M}_{n}$. The knotset operation is defined as $\mathcal{K}\ast\mathcal{M}=\mathcal{K}_{0}\ast\mathcal{K}_{1}\ast\cdots\ast\mathcal{K}_{n-1}\ast\mathcal{M}$ where $\mathcal{K}_{p}=(p,R_{p},K_{p})$ for $p\in[0\,{:}\,n-1]$.
@@
We refer to $\mathcal{M}^{\ast}=\mathcal{K}\ast\mathcal{M}$ for a knotset $\mathcal{K}$ as a knot-model and provide the form of $\mathcal{M}^{\ast}$ in Proposition 3.9. The proof is trivial due to the descending time-ordering of knot applications specified in Definition 3.8. The knotset operator also inherits right-associativity from knots.
@@
Consider a knotset $\mathcal{K}=(M_{0:n-1},K_{0:n-1})$ where $K_{p}=\mathrm{Id}$ for $p\in[0\,{:}\,n-1]$ applied to model $\mathcal{M}=(M_{0:n},G_{0:n})$. The resulting model is $\mathcal{K}\ast\mathcal{M}=\mathcal{M}$.
@@
Though knotsets can change the Feynman–Kac model they are applied to, some quantities remain unchanged, whilst others can be expressed in terms of the original model. We demonstrate these invariances and equivalences now. In order to distinguish quantities related to the original model, $\mathcal{M}$, and to the relevant knot-model, $\mathcal{M}^{\ast}$, we embellish the latter with the same superscript. For example, the terminal measure of $\mathcal{M}$ will be denoted by $\gamma_{n}$, whilst we will use $\gamma_{n}^{\ast}$ for $\mathcal{M}^{\ast}$.
@@ 4 Variance reduction and ordering from knots
Having established the equivalence of all terminal marginal measures in Proposition 3.13, we now consider the asymptotic variances of particle approximations to these quantities. Our main result is given in Theorem 4.1 where we state that applying knots to a Feynman–Kac model reduces the variance of particle approximations for all relevant functions. We will denote the predictive and updated asymptotic variance of the knot-model $\mathcal{M}^{\ast}$ by $\sigma^{2}_{\ast}$ and $\hat{\sigma}^{2}_{\ast}$ respectively.
@@
Note that the variance reduction excludes a contribution from time $n$ due to the absence of a knot at the terminal time. We can also define a terminal time $(n,R,K)$-knot analogously to the knots discussed thus far. However, such a terminal knot will only guarantee a variance reduction of the normalising constant estimate, $\hat{\gamma}_{n}(1)$. We discuss terminal knots and how to achieve variance reductions for specific test functions in Section 6.
@@
The differences in the asymptotic variances stated in Corollary 4.2 are straightforward to derive using the quantitative result in Theorem 4.1 so we suppress them here.